# 2-Cats: 2D Copula Approximating Transforms

### v3

@@ Appendix A Sampling from 2-Cats
Appendix B A Mixture of 2-Cats is a Universal Copula Approximator
With sufficient components, any other CDF (or density function) may be approximated by a mixture of other CDFs [Nguyen et al., 2020]. As a consequence, a mixture of 2-Cats of the form:
@@ Appendix B A Mixture of 2-Cats is a Universal Copula Approximator
This is similar to the Mixture Density NN [Bishop, 1994]. Over the last few years, mixture of Copulas approaches has been gaining traction in several fields [Abraj et al., 2022, Sohrabian, 2021, Zhang and Shi, 2017]; here, we are proposing an NN variation. By definition, this is a valid Copula.
Appendix C On the Integral of Neural Networks
Let $\hat{f}_{\bm{\theta}}(x)$ be some NN parametrized by $\bm{\theta}$. Also, let $f(x)$ represent our training dataset. We also make the reasonable assumption our dataset comes from some real function, $f^{\ast}(x)$, sampled under a symmetric additive noise, i.e., $f(x)=f^{\ast}(x)+\eta$, with an expected value equal to zero: $\mathbb{E}[\eta]=0$. Gaussian noise is such a case.
If a NN is a universal approximator Cybenko [1989], Hornik et al. [1989], Lu et al. [2021], Cai [2023], Lu et al. [2017] and it is trained on the dataset $f(x)$, we have that:
$\lVert\hat{f}_{\bm{\theta}}(x)-f(x)\rVert_{p}<\epsilon_{1}$
Most UAT proofs are valid for any $p$ norm ($p$ norms bound one another), and assuming the 1 norm, we reach: $\hat{f}_{\bm{\theta}}(x)=f(x)\pm\epsilon$ for some constant $\epsilon$. As a consequence, we shall reach the following relationship:
$\hat{f}_{\bm{\theta}}(x)=f^{\ast}(x)+\eta\pm\epsilon_{1}$
Given that integration is a linear operator, we can show that:
$E_{\eta}[\int\hat{f}_{\bm{\theta}}(x)]=\int f^{\ast}(x)\pm\epsilon_{2}.$
$g(x)$ $=\int\hat{f}{{}_{\theta}}(x)\,dx$
[…]
$=\int E_{\eta}[f^{\ast}(x)]+E_{\eta}[\eta]\pm E_{\eta}[\epsilon_{1}]dx$
$=\int E_{\eta}[f^{\ast}(x)]\pm E_{\eta}[\epsilon_{1}]dx$
$=\int f^{\ast}(x)\pm\epsilon_{1}dx$
[…]
where $\epsilon_{2}$ is a constant as $\epsilon_{1}$ is also a constant. However, this constant will grow depending on the integrating interval. Controlling for $\epsilon_{2}$ will depend on the dataset’s quality and the training procedure.
A similar result is discussed in Liu [2023].
Appendix D A Flexible Variation of the Model
@@
    Let $m_{\bm{\theta}}:\mathbb{I}^{2}\mapsto\mathbb{R}+$ be an MLP outputs positive numbers. We achieve this by employing Elu plus one activation in every layer as in Wehenkel and Louppe [2019].
@@
This model meets P3, but not P1 nor P2. As already stated in the introduction, the fact that $t_{v}(u)$ and $t_{u}(v)$ are monotonic, one-to-one, guarantees that $G(t_{v}(u),t_{u}(v))$ defines a valid probability transform to a new CDF. A major issue with this approach is that we have no guarantee that the model’s derivatives are conditional cumulative functions (first derivatives) or density functions (second derivatives). That’s why we call it Flexible (FLEX).
D.1 Full Results
@@
Appendix E Appendix Baseline Source Code

### v4

@@ Appendix B Sampling from 2-Cats
Appendix C Ablation Study
We now provide an ablation study to understand the impact of our three losses. In this study, our model was trained without the Lagrangian terms of Property P3.
Table 3: Ablation Study
BOSTON
$L^{C}$ $L^{dC}$ $L^{c}$
Only $L^{c}$ for training 0.115 0.033 -0.448
$L^{c}$ and $L^{dC}$ for training 0.107 0.035 -0.454
$L^{c}$ and $L^{C}$ for training 0.103 0.023 -0.236
All Three Losses (paper) 0.107 0.020 -0.630
INCT-MSOFT
$L^{C}$ $L^{dC}$ $L^{c}$
Only $L^{c}$ for training 0.131 0.007 -0.327
$L^{c}$ and $L^{dC}$ for training 0.141 0.004 -0.314
$L^{c}$ and $L^{C}$ for training 0.137 0.007 -0.320
All Three Losses (paper) 0.141 0.018 -0.402
GOOG-FB
$L^{C}$ $L^{dC}$ $L^{c}$
Only $L^{c}$ for training 0.166 6.88 -3.324
$L^{c}$ and $L^{dC}$ for training 0.133 0.066 -2.163
$L^{c}$ and $L^{C}$ for training 0.195 2.50 -3.201
All Three Losses (paper) 0.136 0.087 -2.881
The three tables (see Table 3) of this section below present the values for $L^{C}$ (the squared error of the cumulative function C), $L^{dC}$ (the squared error for the first derivatives of C), and $L^{c}$ (the copula likelihood). We present one table for each real-world dataset.
$L^{c}$ is the score of interest when comparing models and the one we report in our paper. However, as discussed in Appendix A of our manuscript, when an NN approximates one aspect of a function (here being the Copula density $L^{c}$), the NN may miss other elements, such as the integral and derivatives of the function. This is the reason why we now present results for all three metrics. The metrics are in the columns of the tables. Each row presents a different 2-Cats approach using: (1) only $L^{c}$ for training (the metric of most interest); (2) $L^{c}$ and $L^{dC}$; $L^{c}$ and $L^{C}$; and, (3) all three losses (as in the main text).
From the tables, we can see that every loss plays a role when training the model. When using all three losses, gains/ties are achieved in 5 out of 9 cases (bold). When ignoring such terms, large losses (italic) will also arise
Appendix D Conjecture: A Mixture of 2-Cats is a Universal Copula Approximator
With sufficient components, any other CDF (or density function) may be approximated by a mixture of other CDFs [52]. As a consequence, a mixture of 2-Cats of the form:
@@ Appendix D Conjecture: A Mixture of 2-Cats is a Universal Copula Approximator
This is similar to the Mixture Density NN [9]. Over the last few years, mixture of Copulas approaches has been gaining traction in several fields [2, 64, 73]; here, we are proposing an NN variation. By definition, this is a valid Copula.
Appendix E Non Scaled Synthetic Data
Table 4: Results for non [0, 1] Scaled Data
Gaussian ($\rho$) Clayton ($\theta$) Frank/Joe ($\theta$) 0.1 0.5 0.9 1 5 10 1 5 10 par $2.91\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.09$ $2.65\pm 0.09$ $1.89\pm 0.11$ $1.39\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.10\pm 0.09$ bern $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.09$ $2.66\pm 0.08$ $2.06\pm 0.09$ $1.76\pm 0.08$ $2.78\pm 0.08$ $2.58\pm 0.09$ $2.15\pm 0.08$ pbern $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.10$ $2.67\pm 0.09$ $2.12\pm 0.10$ $1.92\pm 0.08$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.13\pm 0.08$ pspl1 $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.11$ $2.67\pm 0.09$ $2.14\pm 0.18$ $1.61\pm 0.13$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.10\pm 0.09$ pspl2 $2.91\pm 0.09$ $2.69\pm 0.09$ $2.13\pm 0.10$ $2.67\pm 0.09$ $2.02\pm 0.11$ $1.62\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.11\pm 0.09$ TTL0 $2.92\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.09$ $2.65\pm 0.08$ $2.00\pm 0.21$ $1.46\pm 0.09$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.09$ TLL1 $2.92\pm 0.09$ $2.70\pm 0.10$ $2.09\pm 0.09$ $2.65\pm 0.09$ $1.98\pm 0.21$ $1.43\pm 0.10$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.17\pm 0.20$ TLL2 $2.91\pm 0.09$ $2.75\pm 0.20$ $2.08\pm 0.09$ $2.65\pm 0.09$ $2.03\pm 0.25$ $1.41\pm 0.10$ $2.79\pm 0.09$ $2.57\pm 0.09$ $2.11\pm 0.09$ TLL2nn $2.91\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.10$ $2.65\pm 0.09$ $1.93\pm 0.11$ $1.42\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.09$ MR $2.91\pm 0.09$ $2.70\pm 0.09$ $2.16\pm 0.10$ $2.68\pm 0.08$ $2.01\pm 0.11$ $1.54\pm 0.11$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.11\pm 0.08$ probit $2.91\pm 0.09$ $2.69\pm 0.09$ $2.11\pm 0.09$ $2.66\pm 0.08$ $2.05\pm 0.20$ $1.50\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.08$ 2-Cats-L $\mathbf{2.95\pm 0.09}$ $\mathbf{2.79\pm 0.09}$ $1.91\pm 0.10$ $2.67\pm 0.08$ $2.01\pm 0.09$ $0.97\pm 0.10$ $\mathbf{2.90\pm 0.09}$ $\mathbf{2.73\pm 0.10}$ $2.13\pm 0.10$
Table 4 presents results by scaling the input data as is done in Deep Learning methods. The table presents the average negative log-likelihood and the 95% confidence interval. Overall, statistical ties are present throughout the table.
Parametric methods are expected to overperform in this table, as the data was simulated from the model’s dataset. Nevertheless, the statistical ties with such methods show how 2-Cats is able to capture this behavior even if it is a family-free approach. Some cases where 2-Cats outperformed the baselines are shown in bold.
Appendix F A Flexible Variation of the Model
@@
    Let $m_{\bm{\theta}}:\mathbb{I}^{2}\mapsto\mathbb{R}+$ be an MLP outputs positive numbers. We achieve this by employing Elu plus one activation in every layer as in [71].
@@
This model meets P2, but not P1 nor P3. As already stated in the introduction, the fact that $t_{v}(u)$ and $t_{u}(v)$ are monotonic, one-to-one, guarantees that $G(t_{v}(u),t_{u}(v))$ defines a valid probability transform to a new CDF. A major issue with this approach is that we have no guarantee that the model’s derivatives are conditional cumulative functions (first derivatives) or density functions (second derivatives). That’s why we call it Flexible (FLEX).
F.1 Full Results
@@
Appendix G Appendix Baseline Source Code
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract discusses the model (Section 1); the properties of the model and how they are met (Section 3 and Appendixes); and empirical results (Section 5 and Appendixes).
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Section 3 discusses how some properties are approximated.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: Section 3 presents our formal proofs. Appendix A complements it.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Source code is already released in an anonymous link. Different folders present notebooks for our results.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Source code has already been released. It contains the synthetic data. Real data comes from the literature (also downloaded by our source code).
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Section 5 presents our data splits and random seeds. Our code was also released to verify this. Our model was trained with fixed hyperparameters mentioned in Sections 3 and 5. No tuning was necessary as this was sufficient to overperform baseline methods. The ADAM optimizer is used.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: Our main result tables that compare with baselines present bootstrap confidence intervals.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Section 5 and Appendix F present how baselines were executed and links to source codes. Our code was executed on Google Colab Pro. Our focus on this work was not on scalability, but we do point out that 2-Cats trains in minutes to hours.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: Our research does not deal with human subjects or sensitive data, nor does it have any copyright issues. We evaluate our research on public non-sensitive and synthetic datasets.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. 11.
    Safeguards
  41. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  42. Answer: [N/A]
  43. 12.
    Licenses for existing assets
  44. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  45. Answer: [Yes]
  46. Justification: We only make use of public data and source code for baselines. Our code is already released with am open source license.
  47. 13.
    New Assets
  48. Question: Are new assets introduced in the paper well documented, and is the documentation provided alongside the assets?
  49. Answer: [Yes]
  50. Justification: See the previous question.
  51. 14.
    Crowdsourcing and Research with Human Subjects
  52. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  53. Answer: [N/A]
  54. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  55. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  56. Answer: [N/A]